# FRIDA to the Rescue! Analyzing Synthetic Data Effectiveness in Object-Based Common Sense Reasoning for Disaster Response

### v1

@@ 3.3 Evaluation
We implement 2 different evaluations to compare FRIDA output to the gold-standard evaluation. We first evaluate with an Exact Match (EM) comparison between the gold standard and FRIDA’s output. This is a strict metric, as formatting mistakes (i.e. not predicting the letter corresponding to the answer, adding punctuation) lead to an incorrect mark. We do provide system instructions to help the FRIDA models respond in the correct format, but all prompts are zero-shot. We feel this is a reasonable metric since the instructions are multiple choice, so the model sees the full answer in proper format before responding.
To get a better idea of how close the answers are from a semantic perspective, we also used SemScore Aynetdinov and Akbik (2024); Geronimo and Lera (2024). SemScore is a scoring metric that uses cosine similarity to compare the a tokenizer’s embedding vectors of the gold standard and FRIDA responses. Aynetdinov and Akbik (2024) compared human rankings of 252 LLM instruction responses collected by Wang et al. (2022), and found that SemScore most closely resembled human rankings when compared to other automatic semantic evaluations. Additionally, SemScore was tested with a variety of general instruction based models Aynetdinov and Akbik (2024), making it the best silver-standard metric for understanding how semantically close the FRIDA models were to the gold standard.
@@ 3.4 Ablation Study
As we developed the templates used to generate seed instructions, we wanted to explore which types of data improved model performance the most. To better understand the effectiveness of our data, we ran an ablation study where we fine-tuned all base models on subsets of the synthetic fine-tuning data, which can be seen in Table 2.
We made an ablated model for each category, where each model is fine-tuned only on the synthetic data generated by templates in said category. For example, the Relative Sizes ablation model is trained on data generated from 4 templates testing size, weight, objects fitting in containers, and objects changing state, respectively. We refer to these ablated models as ablated-FRIDA (or aFRIDA) models.
The resulting name for a FRIDA model trained only on data from the Relative Size category would thus be, “aFRIDA 3B: relative sizes” where the “3B” represents the number of parameters in the base model and “relative sizes” refers to the subset of data used for fine-tuning (see Appendix Table 4 for data categories). Since we fine-tuned a LLaMa and Mistral model which both had 8 Billion parameters, we the fine-tuned LLaMa model was named aFRIDA 8B, and the fine-tuned Mistral model was named MaFRIDA 8B (Minstral ablated FRIDA, 8B). The ablated models were tuned with the same hyper-parameters and hardware as the FRIDA models trained on the entire synthetic dataset.
A model suite for a given base model contains FRIDA, trained on the full dataset, as well as 8 aFRIDA models trained on the categorical subsets of the data: relative sizes and state, object function, object differences, specialized equipment, objects causing harm, non-function object facts, earthquake knowledge, and instruction understanding. Examples of data for each category can be found in 4 in the appendix.
@@ 4 Results
Base Models EM Accuracy
LLaMa 3.2 1B Instruct 0.23 LLaMa 3.2 3B Instruct 0.18 LLaMa 3.1 8B Instruct 0.53 Ministral 8B Instruct 0.65
Fine-tuned FRIDA Models EM Accuracy
FRIDA 1B 0.22 FRIDA 3B 0.51 LLaMa FRIDA 8B 0.56 Ministral FRIDA 8B 0.73
aFRIDA 1B: relative sizes 0.34 aFRIDA 3B: relative sizes 0.55 aFRIDA 8B: relative sizes 0.60 MaFRIDA: relative sizes 8B 0.75
aFRIDA 1B: object functions 0.28 aFRIDA 3B: object functions 0.51 aFRIDA 8B: object functions 0.65 MaFRIDA: 8B object functions 0.71
aFRIDA 1B: object differences 0.19 aFRIDA 3B: object differences 0.48 aFRIDA 8B: object differences 0.55 MaFRIDA: object differences 0.66
aFRIDA 1B: objects causing harm 0.27 aFRIDA 3B: objects causing harm 0.45 aFRIDA 8B: objects causing harm 0.56 MaFRIDA: objects causing harm 0.66
aFRIDA 1B: specialized equipment 0.1 aFRIDA 3B: specialized equipment 0.41 aFRIDA 8B: specialized equipment 0.56 MaFRIDA: specialized equipment 0.68
aFRIDA 1B: non-functional obj facts 0.15 aFRIDA 3B: non-functional obj facts 0.49 aFRIDA 8B: non-functional obj facts 0.61 MaFRIDA non-functional obj facts 8B 0.65
aFRIDA 1B: earthquake knowledge 0.25 aFRIDA 3B: earthquake knowledge 0.33 aFRIDA 8B: earthquake knowledge 0.59 MaFRIDA 8B: earthquake knowledge 0.59
aFRIDA 1B: instruction understanding 0.11 aFRIDA 3B: instruction understanding 0.45 aFRIDA 8B: instruction understanding 0.11 MaFRIDA 8B: instruction understanding 0.03
@@
Table 3: The Exact Match Accuracy Scores for the base, FRIDA, and aFRIDA models. While the larger models trained on all data do improve over the base model, the best performance comes from models trained on far smaller subsets of the synthetic dataset involving comparing objects by their physical state or by their functions.
@@
Figure 3: SemScores (embedding-vector cosine similarity scores) for the M-FRIDA 8B and FRIDA 8B models, as well as their corresponding ablation and base models. These heatmaps show that across all models, performance is best in evaluation data corresponding to general common sense (object fuctions, differences) and worst in evaluation data corresponding to specialized object knowledge (earthquake, non-functional object facts). Interestingly, there is little correlation between evaluation category and ablation models trained on synthetic data based on that category.
As seen in Table 3, all FRIDA models larger than 1 Billion parameters scored better in 0-shot exact match scoring than their base models. Surprisingly, the aFRIDA models for the Relative Size and Object Functions categories outperformed their corresponding FRIDA models and pre-trained baselines. The top scoring ablation models consistently outperformed their full FRIDA models by 4-9 points. Minstral-FRIDA 8B not only outperformed LLaMa-FRIDA 8B; it and MaFRIDA 8B: relative sizes also outperformed Gemini-1.5-flash’s Exact Match score of 0.725 in a 0 shot setting.
We also assessed each model’s capability on each subset of templates within the evaluation dataset, and we show the Semscore (cosine similarity) results for the two best fine-tuned models in Figure 3. Overall, models fine-tuned only on objects’ basic physical characteristics and functionality data performed more strongly across all categories. This was despite these categories being the “easiest” in our evaluation, receiving the highest exact match scores and SemScores across all models. It is also clear from both heatmaps that the most difficult topic is equipment-related questions.
Another observation from Figure 3 is that the base models get more consistent scores across the board, while the ablation models are more varied across categories. This is especially true for MaFRIDA 8B, where fine-tuning caused it to perform worse than the base model when measuring by SemScore. Additionally, the highest subset accuracy scores for the aFRIDA models were not always for the data on which they were fine tuned. An example of this is that the model “MaFRIDA 8B: relative size” scored lower on the relative size evaluation templates than the model “MaFRIDA 8B: specialized equipment”, as seen in the leftmost heatmap for M-FRIDA 8B in Figure 3.
5 Discussion
We were surprised by the level of improvement the FRIDA pipeline imbued, especially when our sample for data informativeness was judged to be less than ideal by the authors. We were especially surprised that the aFRIDA relative size and object function models outperformed all other models across the board. It’s possible that clarifying the basic properties and affordances of objects provided a better basis for the model to have better reasoning. Another contributing factor is likely that our synthetic data on specific objects and tasks tended to be longer, more diverse, and less informative according to our dataset analysis. The sample size of these more specific synthetic data may have been too small for major improvement in those areas.
An additional unexpected observation was that the Instruction Understanding ablation models performed so poorly in the exact match metric, but comparably to the other aFRIDA models with SemScore. We believe this is because the templates for instruction understanding included punctuation in the answer choices, but none of the other templates did. Thus the models fine-tuned only of instruction understanding would add punctuation and miss the Exact Match, reinforcing the need for multiple types of evaluation on this task.
Overall, the FRIDA pipeline is effective at improving small LLMs on its data. Another positive is that they are very lightweight, with comparable performance to a much larger Gemini model. The data relating to objects function and physicality seem to improve model performances on both the Exact Match and the SemScore metrics. It is clear, however, that FRIDA models could improve with different data distributions more heavily favoring object size and functions (which helped it improve) and specialized equipment (which all models struggled on). The base models clearly have enough access to this disaster to correctly answer the more fact-based templates regarding disaster response, while physical common sense-related questions help to improve real-world knowledge for practical interactions.
6 Related Work
6.1 LLMs Reasoning about the World
There a wide variety of methods for leveraging LLMs for reasoning in a physical environment based on Chain of Thought prompting Wei et al. (2022b). These include variants like re-prompting Raman et al. (2022), which prompts the LLM to regenerate a plan if certain criteria aren’t met at each steps, or Tree of Thought Yao et al. (2023), which generates a tree of potential steps and evaluates each potential path via either a breadth-first or depth-first search.
There are also methods that take allow the LLM to take in environmental feedback in response to its output. For Inner-Monologue Huang et al. (2023), the LLM is given the option to ask for more scene descriptors from a human handler, which it then incorporates into its prompts, improving task completion and decreasing hallucination. Another example is SayPlan Rana et al. (2023), which uses 3D scene plans to iterate on proposed strategies until an effective path is discovered. Xie and Zou (2024) get feedback from LLMs themselves by using a wide variety of LLM agents to do various sub-tasks for planning, including generating a general outline, using external tools to gain information, and evaluating which plan is best.
One resource for improving LLM understanding object affordances specifically is Adak et al. (2024), who curate a dataset of naturally occurring sentences and corresponding images, then transform them into inference, probing, and text and visual masking tasks. They further prove that even Visual Language Models (VLMs) do not have straightforward understandings of affordances, but few-shot fine-tuning improved LLM and VLM performance on identifying object affordances.
6.2 Disaster Work and Natural Language Processing
Godinho (2024) completed a systematic search and analysis of over 100 peer-reviewed papers relating to Natural Language Processing (NLP) tools being applied to disasters. 85 of the 107 papers found were for analyzing social media, with 67 of the papers analyzing twitter data specifically. Over half of the total papers had a sentiment analysis component to their work, and the 2nd and 3rd most common tools used were text classification and information extraction. 87.8% of papers focused on natural disasters, with only 3.7% being solely about man-made disasters (the rest pertained to both) Godinho (2024). Godinho (2024) and Wang et al. (2024) together showcase that while there is much research on LLMs as agents and much research on NLP analysis of disasters, there is not much overlap in these spaces.
7 Future Work
@@
Even with strong performance, we still feel there are several ways we can further improve the FRIDA pipeline. Firstly, we want to improve our prompting for synthetic data to make them less trivial. We want to refine our common sense related templates and make sub-templates with different phrasing to make our synthetic data more reflective of real world natural language. We also hope implementing the strategies in other work Ge et al. (2024); Ding et al. (2023); Mukherjee et al. (2023) for diversifying synthetic data will improve quality and make data generation more time-efficient. Finally, we plan to test the pipeline on a variety of specific disasters with disaster experts who can provide feedback on the feasibility of our process.
@@ 8 Conclusion
We introduce a pipeline to create expert-knowledge-based synthetic data that is then used for fine-tuning to create FRIDA models. We found our pipeline substantially improved performance over 3 different instruction-tuned LLaMa models and 1 instruction-tuned Ministral model. We then performed an ablation study and found that data generated from templates based in physical common sense reasoning about objects improved performance most; ablation models trained on those data scored higher than FRIDA models trained on all synthetically generated data. This pipeline is an important step in understanding and improving LLM object reasoning for practical use.
@@ 9 Limitations, Risks, and Ethics
One limitation is that we train and evaluate on template-generated data rather than naturally occurring language; there could be linguistic or stylistic differences between template-generated data and naturally occurring instructions. Though our approach still relies on access to expert input and non-trivial computational power for fine-tuning to counter these shortcomings, we outline solutions in Section 7 which we believe are ripe avenues for future work.
@@ Appendix A Categories and Descriptions
Relative Sizes Biggest Object, Heaviest Object, Relative Fit Which of these objects is the lightest? outlet, broom, pail, orange, screen 20
@@
Appendix B Synthetic Data Histograms by Category
Appendix C Fine Tuning Specifics
@@
Appendix D Synthetic Data Generation Prompting
@@
Appendix E Licenses
We used TRL von Werra et al. (2020) under the Apache License. SemScore Geronimo and Lera (2024) implements the MIT license, and the LLaMa models were used after author agreement to the LLaMa 3.1 and 3.2 Community License Agreement Team (2024b). Ministral 8B Instruct was used under the Mistral Research License Jiang et al. (2023). We used the Disagree github to calculate inter-annotator agreement scores under the MIT license Price and Mensio (2022). We will release our code, datasets, and FRIDA models under a open-source license which will be chosen upon conference acceptance. Team (2024c).

### v2

@@ 3.4 Evaluation
Although we leverage multiple choice questions and answers for evaluation, we required a less rigid method than exact match so that formatting errors (e.g. writing “A” instead of “A)”, or forgetting punctuation) would have less impact. Thus, we used SemScore Aynetdinov and Akbik (2024); Geronimo and Lera (2024), which is a scoring metric that uses cosine similarity to compare a model’s embedding vectors of the gold standard and FRIDA responses.
@@ 3.5 Ablation Study
To better understand the effectiveness of the types of physical reasoning represented in our synthetic data, we ran an ablation study where we fine-tuned our base model on subsets of the synthetic fine-tuning data, which can be seen in Table 2. We made an ablated model for each category of data, where each model is fine-tuned only on the synthetic data generated by templates in said category. For example, the “Relative Sizes and Shapes” ablation model is trained on data generated from 4 templates testing size, weight, objects fitting in containers, and objects changing state. We refer to these ablated models as ablated-FRIDA (or aFRIDA) models.
The resulting name for a FRIDA model trained only on data from the Relative Sizes and Shapes category would thus be, “aFRIDA: relative sizes and shapes”, where “relative sizes and shapes” refers to the subset of data used (see Appendix Table 4 for data categories). The ablated models were tuned with the same hyper-parameters and hardware as the full FRIDA model.
A model suite for a given base model contains FRIDA, trained on the full dataset, as well as 8 aFRIDA models trained on the categorical subsets of the data: relative sizes and shapes, object function, object differences, specialized equipment, objects causing harm, non-function object facts, earthquake knowledge, and instruction understanding. Examples of data for each category can be found in Table 4 in the appendix.
@@ 4 Results
Model SemScore Accuracy (%)
Ministral 8B Instruct 93.5
FRIDA 94.6
aFRIDA Accuracy (%),
Fine tuning Data Subset all evaluation data
relative sizes and shapes 95.0
object functions 94.7
object differences 93.4
objects causing harm 93.3
specialized equipment 93.8
non-functional obj facts 93.2
earthquake knowledge 91.7
instruction understanding 85.0
@@
Table 3: The SemScore Accuracy on all evaluation data for the base model Ministral 8B Instruct, the fine-tuned FRIDA model trained on all synthetic data, and the fine-tuned models trained on ablated subsets of the synthetic data (aFRIDA). The FRIDA model trained on all data improved performance over its corresponding base model. The best overall performance came from the aFRIDA model trained on a subset of the synthetic dataset involving comparing objects by their physical state.
As seen in Table 3, the Ministral 8B FRIDA model had a higher SemScore Aacuracy than its base model. However, the aFRIDA models for the “Relative Size and Shape” and “Object Functions” categories outperformed both the unablated FRIDA model and the base model. These models also outperformed Gemini-1.5-flash’s SemScore of 93.9 in a zero shot setting.
We assessed each model’s capability on each type of reasoning tested in the evaluation dataset. To show the overall trend across models, we present the SemScore results for the FRIDA and aFRIDA models in Figure 4. Overall, when observing model performance in the Figure 4’s columns, models fine-tuned only on objects’ basic size and shape characteristics or only on object functionality performed more strongly across most evaluation categories. This was despite these synthetic data covering straightforward physical semantics that don’t require any highly specific knowledge or creativity like the “specialized equipment” or “objects causing harm” categories. These models also had the strongest performance with far less training data than the full FRIDA model (see Table 2).
Looking at evaluation data types represented in the rows, it is clear that the more difficult evaluations are “specialized equipment”, the category querying about the specialized objects used in earthquake search and rescue, and “earthquake”, the category evaluating scientific knowledge about earthquakes. Both of these evaluations are highly specific and technical. The easier evaluation categories are “object functions” and “differences”, which pertain to understanding the basic semantics of objects’ abilities and the differences between objects, respectively.
@@
Figure 4: SemScores (embedding-vector cosine similarity scores) for the FRIDA suite for each type of evaluation. Across all models, performance is better in evaluation data corresponding to physical common sense (object functions, differences) and worse in evaluation data corresponding to specialized object knowledge (earthquake, specialized equipment).
Another key observation from Figure 4 can be found by comparing evaluation performance between FRIDA and Ministral 8B. FRIDA has stronger performance than the base model except for the “required equipment”, “earthquake”, and “instruction following” evaluations. This could potentially demonstrate that these data need to be generated differently or that Ministral 8B needs more of them in order to strengthen performance.
5 Discussion & Error Analysis
It is particularly surprising that the “aFRIDA relative size and shape” and the “aFRIDA object function” models outperformed all other models across the board, even though the physical semantics expressed in those fine-tuning data are not complex. We hypothesize that clarifying the basic properties and affordances of objects provided a better basis for the model to have stronger physical reasoning across all categories.
Another surprise was that the “relative sizes and shapes” evaluation subset was a challenge for the FRIDA suite. Although one may think that simpler object properties like its “relative sizes and shapes” might be relatively prevalent in the base models’ pre-training data, it is also plausible that reporting bias in web text leads to under-representation of highly commonplace facts Raji et al. . We hypothesize that this lack of pretraining data is partially why the ablation model trained on “relative shapes and sizes” synthetic data performs so strongly. However, this does not answer why the ablated models trained on data pertaining to other challenging categories in our evaluation, namely “aFRIDA: earthquake” and “aFRIDA: specialized equipment”, did not receive the same overall performance bump.
We suspect that the reason “aFRIDA: earthquake” and “aFRIDA: specialized equipment” did not similarly improve performance is that our synthetic data for the more specific objects and tasks tended to be longer and have lower ROUGE scores. These data therefore had more diversity. The sample size of the Earthquakes and Specialized Equipment synthetic data subsets may have been too small for the model to be effectively biased by fine tuning. Conversely, larger models may have ingested operators’ manuals for specialized equipment, facilitating parroting answers for questions on this topic. We note that our research highlights the general difficulty of analyzing the precise effects of fine-tuning given opaque pre-training data.
Error analysis of both the FRIDA model (fine-tuned using all synthetic data) and the “aFRIDA relative size and shape” model revealed that both models got the same instances and number of the “relative size and shape” evaluation data incorrect. For example:
  1. 1.
    “What is the easiest way to use a camera?”
    A) with the camera plugged in
    B) with the camera unplugged
    Gold: B) with the camera unplugged
    FRIDA: A) with the camera plugged in
The base Ministral model generally gets the same “relative size and shape” evaluation instances incorrect as the FRIDA models. However, it also answers incorrectly for over half of the instances of test items that relate to answering which item is bigger and which item will fit into another item. For example:
  1. 2.
    “Choose the biggest of a given set of objects in terms of your own common sense”
    A) bicycle
    B) chalk
    C) poster
    D) jar
    E) taillight
    Gold: A) bicycle
    Ministral: D) jar
  2. 3.
    “Can chalk fit in a cup?”
    Answer “it can” or “it cannot”
    Gold: it can
    Ministral: it cannot
Thus, we conclude that the fine-tuning contributed to improvement in understanding which items are bigger and which items fit into others in particular. This improvement may translate to improvement in other related categories. Specifically, we also see dramatic improvement over the base model for the “objects causing harm” evaluation data. This could be further boosted by a general understanding of which objects are larger.
Overall, the FRIDA pipeline improves small LLM object reasoning when said models are fine-tuned on more general physical common sense and object reasoning data. The FRIDA suite models are lightweight enough to fit within our constraints, and can even achieve comparable performance to a much larger Gemini model. In comparison to the ablated models, the performance of the full FRIDA model trained on all synthetic data demonstrates that more work needs to be done to improve the synthetic dataset distribution to be ideal for improving FRIDA model performance on reasoning for earthquake search and rescue.
5.1 Future Work
@@
There are several ways we can further improve the FRIDA pipeline. We want to improve our prompting for synthetic data to make them less trivial. We can refine and expand our less technical templates. By adding different phrasing, we hope to make our synthetic data more reflective of real world natural language. We also hope implementing the strategies in other work Ge et al. (2024); Ding et al. (2023); Mukherjee et al. (2023) for diversifying synthetic data will improve generation quality and efficiency. Finally, we plan to test the pipeline on other domains with experts to help us refine our process.
@@ 6 Conclusion
We introduce a pipeline to create expert-in-the-loop-based synthetic data that is then used for fine-tuning to create FRIDA models. We found our pipeline improved performance over our base model. We performed an ablation study and found that data generated from templates based in basic physical common sense reasoning about objects improved performance most; ablated models trained on those data scored higher than FRIDA models trained on all synthetically generated data and higher than Gemini-1.5-flash, the LLM that generated the synthetic data. This pipeline is an important step in understanding and improving LLM object reasoning for practical use. Even if some of our problem constraints are eventually alleviated by technology that facilitates very large models with smaller compute requirements, there will remain problem spaces for which web-based pre-training data simply does not exist. Our research demonstrates an effective pipeline to specialize models fine-tuned on data that is not well-represented in typical web text pre-training data.
@@ 7 Limitations, Risks, and Ethics
One limitation is that we train and evaluate on template-generated data rather than naturally occurring language; there could be linguistic or stylistic differences between template-generated data and naturally occurring instructions. Though our approach still relies on access to expert input and non-trivial computational power for fine-tuning to counter these shortcomings, we outline solutions in Section 5.1 which we believe are ripe avenues for future work.
@@ Appendix A Categories and Descriptions
relative sizes and shapes Biggest Object, Heaviest Object, Relative Fit Which of these objects is the lightest? outlet, broom, pail, orange, screen 20
@@
Appendix B Fine Tuning Specifics
@@
Appendix C Synthetic Data Generation Prompting
@@
Appendix D Licenses
We used TRL von Werra et al. (2020) under the Apache License. SemScore Geronimo and Lera (2024) implements the MIT license, and the LLaMa models were used after author agreement to the LLaMa 3.1 and 3.2 Community License Agreement Team (2024b). Ministral 8B Instruct was used under the Mistral Research License Jiang et al. (2023). We will release our code, datasets, and FRIDA models under a open-source license which will be chosen upon conference acceptance.